OpenStudy (anonymous):

factoring x^6-26x^3-27=

OpenStudy (anonymous):

ok put x^3 = t

OpenStudy (anonymous):

then the equation become t^2 - 26 t -27

OpenStudy (anonymous):

now when we factorise them t^2 + 27t -t -27 = 0 t(t+27) -1(t+27)=0 (t-1)(t+27)=0 : )

OpenStudy (anonymous):

now t = -27, 1 t = x^3 => x^3 = -27 x = -3 and similarly t = 1 => x =1

OpenStudy (anonymous):

t^2 + 27t -t -27 = 0 it is not t^2-26t-t-27

OpenStudy (anonymous):

it not make sense

OpenStudy (anonymous):

it can be t^2-27t+t-27

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

yea sure it can be the u sttd above ....sry i ..got it wrong ...but my method is right ..so u follow tht

OpenStudy (anonymous):

did u get it ...the main step here is x^3=t one ...

OpenStudy (anonymous):

ok dear

OpenStudy (anonymous):

now frm the right equation u will get t = 27 ,-1

OpenStudy (anonymous):

then when u put t= x^3 ...u will get the answers ..

OpenStudy (anonymous):

so it is (t-1)(t-27)=o

OpenStudy (anonymous):

no ... rmmbr the eq t^2 - 27t +t -27 =0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now t(t - 27 ) +1 (t - 27) =0

OpenStudy (anonymous):

this gives u (t +1 )(t - 27) = 0

OpenStudy (anonymous):

which gives t = -1 , 27

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now u solve it for x u know t = x^3 i'm doing one ...root .. u tell me nther one x^3 = 27 x = ((3)^3)^1/3 x=3

OpenStudy (anonymous):

so it is (x^3+1)(x^3+3^3)

OpenStudy (anonymous):

op x^3-3^3

OpenStudy (anonymous):

yea sry typo done by me .... u r right

OpenStudy (anonymous):

so you have formular a^3 + b^3

OpenStudy (anonymous):

and a^3-b^3

OpenStudy (anonymous):

but here u dont need it ... wt u can do is (x^3 + 1)(x^3 -3^3) = 0 the u divide both sides by (x^3 - 3^3) then u get (x^3 + 1 ) = 0 . 1 /(x^3 -3^3) here we know nethng multiplied by zero is zero so x^3 = -1 x = -1 u get one root by this ...solve the sam way fr another one

OpenStudy (anonymous):

though (a^3 - b^3) = (a-b)(a^2 + b^2 +ab)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u can use tht one formula too a^3 - b^3 but it will make it unnecessary lengthy

OpenStudy (anonymous):

its result is (x+1)x^2-x+1)(x-3)(x^2+3x+9)

OpenStudy (anonymous):

yea thts another way of saying thts all i thought u have to solve fr x .....i didn't u were to make factors ...yea ur solution is CORRECT !

OpenStudy (anonymous):

thank for your help too

OpenStudy (anonymous):

: )